What is the value of△Go at Amn 25 °C for the reaction -98.8 kJTmo 1-1 141.5 Jmol-1.K-1 55. below? 2Fe(s) + 3H20(g) → FeO(s) + 3H2(g) (A) 4.21x10 kJ-mol1 (B)-56.6 kJ-mol-1 (C)-95.3 kJ-mol-1 (D) -141 kJ-mol- AOr (kJ-mol-1) | What is the value of△Gof of s02(g) using the information to。(g) 0.0 the right and below? 2SO2(g) + O2(g) → 2SOg(g) Δ0%,--142.0 kJ-mol-1 56. SO(g) -371.1 (A) -884.2 kJ-mol-1 (C) 442.1 kJ-mol1 (B) -600.2 kJ-mol-1 (D) -300.1 kJ-mol-1 57. What can be concluded about the sign of AH and AS from this graph? 61 150 100 D -50 -100 150 100 200 300 400 500 600 700 (A) ΔH < 0, AS < 0 (C) ΔH < 0, AS > 0 Temperature (K) (B) (D) ΔH > 0, AS < 0 ΔH > 0, AS > 0
Answer #1

Q55

dG = dH - T*dS

dG = -98.8 -298*-141.5/1000

dG= -56.633 kJ/mol

choose B

Q56

dG = 2SO 3- (2SO2 + O2)

-142 = 2*-371 - (2*SO2 + 0)

SO2 =( -142 - 2*-371)/-2

dG-SO2 = -300

choose D

Q57

given that dG becomes more possitive as T increases, then

dG =H -T*dS

h must be negative,

as T increases, dG becomes more positive, so dS must be negative

then

dH <0, dS < 0